Find the variation constant and an equation of variation where Y varies directly as X and Y=60 when X=10.
The variation constant is K=??
The equation of variation is Y=??
Please explain how you arrived at this answer. Thanks!

Answer by stanbon(75887)

Find the variation constant and an equation of variation where Y varies directly as X and Y=60 when X=10.
----------------
Direct Variation: y = kx
Find k if y = 60 when x = 10:
60 = k*10
k = 60/10 = 6
---------------------
The variation constant is K=6
-------
The equation of variation is Y=6x
==============
Cheers,
Stan H.